In the triangle abc ,the coordinates of b are (0,0),AB=2,angle ABC=pi/3 and the middle point of BC has the coordinates (2,0).Then the centroid is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
slope of BC = 0, therefore this is the x axis
       Coordinates of C = (4,0)
       Coordinates of A = (0+2cos (π/3), 0+2sin(π/3))
= (1,√3)
Centroid= ( 4+0+1)/3 , (0+√3+0)/3
             =( 5/3, 1/√3 )
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)